OpenStudy (anonymous):

A circle is centered at (−3, 2) and has a radius of 2. Which of the following is the equation for this circle? A. (x + 3)2 + (y − 2)2 = 2 B. (x − 3)2 + (y + 2)2 = 4 C. (x + 3)2 + (y − 2)2 = 4 D. (x − 3)2 + (y + 2)2 = 2

OpenStudy (alexandervonhumboldt2):



OpenStudy (alexandervonhumboldt2):

in your case (x, y)=(-3, 2)

OpenStudy (alexandervonhumboldt2):

substitute your input and get your eqaution

OpenStudy (anonymous):

So, is the answer D?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

your point is (-3, 2) thus (x-(-3))^2+(y-2)^2=...

OpenStudy (alexandervonhumboldt2):

no as well -(-)=3 thus we have (x-(-3))^2+(y-2)^2=... =>(x+3)^2.......

OpenStudy (alexandervonhumboldt2):

woops

OpenStudy (alexandervonhumboldt2):

-(-)=+ i meant

OpenStudy (anonymous):

A?

OpenStudy (alexandervonhumboldt2):

nearly: left part is correct, but the right one is incorrect: because the right sides = r^2, not r